# Query Performance Prediction using Relevance Judgments Generated by Large Language Models

### v1

@@ 1. Introduction
Limitations. QPP methods can be applied in various domains and scenarios (Faggioli et al., 2023b); we are usually concerned with the predicted retrieval quality w.r.t. various IR measures across different scenarios, e.g., our emphasis might be on precision (Faggioli et al., 2023d, 2021a) for conversational search and on recall for legal search (Tomlinson et al., 2007). However, existing QPP approaches typically predict only a single real-valued score that indicates the retrieval quality for a query (Ganguly et al., 2022) and do not require the predicted score to approximate a specific IR evaluation measure (Arabzadeh et al., 2023; Singh et al., 2023; Tao and Wu, 2014; Shtok et al., 2012; Zhou and Croft, 2007). This results in two key limitations: (i) While the predicted score has been shown to correlate with some IR evaluation metrics (Datta et al., 2022c; Ganguly et al., 2022), relying on a single value to represent different IR evaluation measures leads to a “one size fits all” issue, which is problematic because the literature shows that some IR metrics do not correlate well and the agreement varies across scenarios and queries (Gupta et al., 2019; Jones et al., 2015), Although some studies train regression-based QPP models to predict a specific IR evaluation measure (Khodabakhsh and Bagheri, 2023; Chen et al., 2022; Datta et al., 2022a; Arabzadeh et al., 2021b; Hashemi et al., 2019), they require training separate models to predict different measures, leading to lots of storage and running costs. (ii) A single-score prediction limits the interpretability of QPP; it is insufficient to explain QPP outputs or to analyze and fix inaccurate QPP results based solely on a single score. We argue that more in-depth and interpretable insights into QPP outputs are required.
@@
A new QPP framework. We propose a QPP framework using automatically generated relevance judgments (QPP-GenRE), in which we decompose QPP into independent subtasks of automatically judging the relevance of each item in a ranked list to a given query. QPP-GenRE comes with various advantages: (i) it allows us to directly predict any desired IR evaluation measure at no additional cost, using generated relevance judgments as pseudo-labels; and (ii) the generated relevance judgments provide an explanation beyond simply gauging how difficult or easy a query is by offering information about why the query is predicted as difficult or easy; moreover, we can translate the “QPP errors” into easily observable “relevance judgment errors,” e.g., false positives or negatives, informing potential ways of improving QPP quality by fixing observed relevance judgment errors.
@@
QPP-GenRE can be integrated with various approaches for judging relevance. Obviously, for QPP-GenRE to excel in QPP, it is crucial to equip QPP-GenRE with an approach capable of accurately generating relevance judgments. Recent studies (Thomas et al., 2023; Faggioli et al., 2023a; MacAvaney and Soldaini, 2023; Gilardi et al., 2023) have shown the potential effectiveness of using LLMs to generate relevance judgments. We build on them to equip QPP-GenRE with LLMs for judging relevance. However, those studies have certain limitations: (Thomas et al., 2023; Faggioli et al., 2023a) prompt commercial LLMs (e.g., ChatGPT, GPT-3.5/4) to generate relevance judgments; commercial LLMs come with limitations like non-reproducibility, non-deterministic outputs and potential data leakage between pretraining and evaluation data, impeding their utility in scientific research (Pradeep et al., 2023b; Zhang et al., 2023b; Pradeep et al., 2023a); although MacAvaney and Soldaini (2023) prompt small-scale open-source language models (e.g., Flan-T5 (Chung et al., 2022) with 3B parameters) for generating relevance judgments, they focus on a setting wherein the model is already given one relevant item for each query, which does not apply to QPP as we typically do not know any relevant item for a query in advance. In this paper, we investigate the possibility of leveraging a leading open-source LLM, LLaMA (Touvron et al., 2023b, a) for generating relevance judgments; LLaMA has shown competitive performance on many natural language processing (NLP) and IR tasks (Touvron et al., 2023b, a). Also, we focus on a more realistic setting wherein we do not know any relevant items for a query in advance.
@@
Challenges. We face two challenges when using QPP-GenRE for QPP: (i) predicting recall-oriented measures requires seeking all relevant items in the corpus for a query; it means that QPP-GenRE needs to judge the entire corpus per query, which is infeasible due to the high computational costs; (ii) our experiments reveal that directly prompting LLaMA in a zero-/few-shot way yields limited effectiveness in judging relevance, resulting in limited QPP quality; this aligns with recent findings indicating limited success in prompting open-source LLMs for specific tasks (Qin et al., 2023).
@@
Solutions. To address the challenges, (i) we devise an approximation strategy to predict recall-oriented IR measures by only judging a few items in the ranked list for a query and using them to estimate the metric, hence avoiding the cost of traversing the entire corpus to find all relevant items for a query; the approximation strategy also enables us to investigate the impact of various judging depths in the ranked list on QPP quality; and (ii) we enhance LLaMA’s ability to generate relevance judgments by training it with parameter-efficient fine-tuning (PEFT) (Dettmers et al., 2023) on human-labeled relevance judgments; unlike previous supervised QPP methods that need to train separate models for predicting different IR evaluation measures, training LLaMA to judge relevance is agnostic to a specific IR metric.
@@
Experiments. Experiments on the TREC 2019–2022 deep learning (TREC-DL) tracks (Craswell et al., 2022, 2021, 2020, 2019) show that QPP-GenRE achieves state-of-the-art QPP performance in estimating the retrieval quality of a lexical (BM25) and a neural ranker (ANCE, Xiong et al., 2021) in terms of both precision- (RR@10) and recall-oriented IR evaluation metrics (nDCG@10). Also, we find that using LLMs to directly model QPP, i.e., asking LLMs to directly generate values of IR evaluation metrics, performs much worse than QPP-GenRE. This observation reveals that QPP-GenRE is a much more effective way of modeling QPP using LLMs.
Furthermore, our experiments demonstrate the effectiveness of our devised approximation strategy in predicting recall-oriented IR measures: QPP-GenRE achieves state-of-the-art QPP performance at the shallow judging depth 10, and further judging up to 100–200 retrieved items in a ranked list can reach saturation. Moreover, we compare QPP-GenRE using fine-tuned LLaMA against few-shot prompted LLaMA, which reveals that better quality in generating relevance judgments translates into better QPP quality. The performance of fine-tuned LLaMA in terms of judging relevance exceeds that of few-shot prompted LLaMA by a large margin, even outperforming that of a commercial LLM (GPT-3.5) (Faggioli et al., 2023a). Touvron et al. (2023b, a) show that there is no data leakage between LLaMA’s pre-training and our evaluation sets. The poor performance of non-fine-tuned LLaMA in judging relevance, further confirms this statement. As an important aside, providing such assurance is infeasible for commercial LLMs. Besides, we analyze QPP errors based on automatically generated relevance judgments, demonstrating QPP-GenRE’s interpretability.
@@
Finally, our computational cost analysis shows that QPP-GenRE shows lower latency than some supervised QPP baselines when predicting multiple measures because multiple measures can be derived from the same set of relevance judgments. QPP-GenRE does show higher latency than those baselines when predicting only one measure. However, given QPP-GenRE’s high effectiveness, it can well fit some knowledge-intensive scenarios where QPP quality is prioritized and users may have a higher tolerance level for latency than users in web search, e.g., legal (Tomlinson et al., 2007) or patent search (Lupu et al., 2013). Plus, QPP can be used to analyze how well a search system performs in offline settings (Faggioli et al., 2023d), where latency is not necessarily an issue.
Contributions. Our main contributions are as follows:
@@
We propose a novel QPP framework using automatically generated relevance judgments (QPP-GenRE), which decomposes QPP into independent subtasks of judging the relevance of each item in a ranked list to the query, and predicts different IR evaluation measures based on generated relevance judgments.
@@
We devise an approximation strategy to predict recall-oriented IR measures, avoiding the cost of traversing the entire corpus to find all relevant items for a query.
@@
We fine-tune a leading open-source LLM, LLaMA, on automatically generating relevance judgments. We release our code and checkpoints at https://github.com/ChuanMeng/QPP-GenRE.
@@
We conduct experiments on four datasets, showing that QPP-GenRE outperforms the state-of-the-art QPP baselines on TREC-DL 19–22 collections in predicting RR@10 and nDCG@10 in terms of Pearson’s and Kendall’s .

### v2

@@ Abstract.
We face two main challenges: (i) excessive computational costs of judging an entire corpus for predicting a metric considering recall, and (ii) limited performance in prompting open-source LLMs in a zero-/few-shot manner. To solve the challenges, we devise an approximation strategy to predict an IR measure considering recall and propose to fine-tune open-source LLMs using human-labeled relevance judgments. Experiments on the TREC 2019–2022 deep learning tracks show that QPP-GenRE achieves state-of-the-art QPP quality for both lexical and neural rankers.
@@ 1. Introduction
QPP methods can be applied in various domains and scenarios (Faggioli et al., 2023b). We are usually concerned with the predicted retrieval quality w.r.t. various IR measures across different scenarios, e.g., our emphasis might be on precision (Faggioli et al., 2023d, 2021a) for conversational search and on recall for legal search (Tomlinson et al., 2007). However, existing QPP approaches typically predict only a single real-valued score that indicates the retrieval quality for a query (Ganguly et al., 2022) and do not require the predicted score to approximate a specific IR evaluation measure (Arabzadeh et al., 2023; Singh et al., 2023; Tao and Wu, 2014; Shtok et al., 2012; Zhou and Croft, 2007). This results in two key limitations: (i) While predicted performance scores have been shown to correlate with some IR evaluation metrics (Datta et al., 2022c; Ganguly et al., 2022), relying on a single value to represent different IR evaluation measures leads to a “one size fits all” approach, which is problematic because the literature shows that some IR metrics do not correlate well and the agreement varies across scenarios and queries (Gupta et al., 2019; Jones et al., 2015). Although some studies train regression-based QPP models to predict a specific IR evaluation measure (Khodabakhsh and Bagheri, 2023; Chen et al., 2022; Datta et al., 2022a; Arabzadeh et al., 2021b; Hashemi et al., 2019), they require training separate models to predict different measures, leading to lots of storage and running costs. (ii) A single-score prediction limits the interpretability of QPP. It is insufficient to explain QPP outputs or to analyze and fix inaccurate QPP results based solely on a single score. We argue that more in-depth and interpretable insights into QPP outputs are required.
A novel QPP framework
@@
We propose a QPP framework using automatically generated relevance judgments (QPP-GenRE), in which we decompose QPP into independent subtasks of automatically predicting the relevance of each item in a ranked list for a given query. QPP-GenRE comes with various advantages: (i) it allows us to directly predict any desired IR evaluation measure at no additional cost, using automatically generated relevance judgments as pseudo-labels; and (ii) the generated relevance judgments provide an explanation beyond simply gauging how difficult or easy a query is by offering information about why the query is predicted as being difficult or easy; moreover, we can translate the “QPP errors” into easily observable “relevance judgment errors,” e.g., false positives or negatives, informing potential ways of improving QPP quality by fixing observed relevance judgment errors.
@@
QPP-GenRE can be integrated with various approaches for judging relevance. Obviously, for QPP-GenRE to excel in QPP, it is crucial to equip QPP-GenRE with an approach capable of accurately generating relevance judgments. Recent studies (Rahmani et al., 2024; Ma et al., 2024; Thomas et al., 2023; Faggioli et al., 2023a; MacAvaney and Soldaini, 2023; Gilardi et al., 2023) have shown the potential effectiveness of using LLMs to generate relevance judgments. We build on them to equip QPP-GenRE with LLMs for judging relevance. However, those studies have certain limitations: several authors have prompted commercial LLMs (e.g., ChatGPT, GPT-3.5/4, GPT-4o) to generate relevance judgments (e.g., Upadhyay et al., 2024b; Rahmani et al., 2024; Upadhyay et al., 2024a; Zhang et al., 2024; Ma et al., 2024; Thomas et al., 2023; Faggioli et al., 2023a); commercial LLMs come with limitations like non-reproducibility, non-deterministic outputs and potential data leakage between pretraining and evaluation data, impeding their value in scientific research (Pradeep et al., 2023b; Zhang et al., 2023c; Pradeep et al., 2023a). Although MacAvaney and Soldaini (2023) prompt small-scale open-source language models (e.g., Flan-T5 (Chung et al., 2022) with 3B parameters) for generating relevance judgments, they focus on a setting wherein the model is already given one relevant item for each query, which does not apply to QPP as we typically do not know any relevant item for a query in advance. In this paper, we focus on the use of open-source LLMs for generating relevance judgments in a realistic setting where we lack prior knowledge of any relevant items for a query. There are only few studies (Zhang et al., 2024; Salemi and Zamani, 2024; Upadhyay et al., 2024a; Khramtsova et al., 2024) attempting to prompt open-source LLMs in this setting.
Challenges
@@
We face two challenges when using QPP-GenRE for QPP: (i) predicting IR metrics considering recall, ideally, entails identifying all relevant items in the entire corpus for a query; however, using an LLM to judge the entire corpus per query is impractical due to the significant computational overhead; (ii) our experiments reveal that directly prompting open-source LLMs in a zero-/few-shot manner yields limited effectiveness in predicting relevance, resulting in limited QPP quality; this aligns with recent findings indicating limited success in prompting open-source LLMs for specific tasks (Qin et al., 2023).
Solutions
@@
To address the challenges listed above, (i) we devise an approximation strategy to predict IR measures considering recall by only judging a few items in the ranked list for a query and using them to estimate the metric, hence avoiding the cost of traversing the entire corpus to identify all relevant items for a query; the approximation strategy also enables us to investigate the impact of various judging depths in the ranked list on QPP quality; and (ii) we enhance an open-source LLM’s ability to generate relevance judgments by training it with parameter-efficient fine-tuning (PEFT) (Dettmers et al., 2023) on human-labeled relevance judgments; unlike previous supervised QPP methods that need to train separate models for predicting different IR evaluation measures, training LLMs to judge relevance is agnostic to a specific IR metric.
Experiments
We integrate QPP-GenRE with a leading open-source LLM, LLaMA (Touvron et al., 2023b, a) for generating relevance judgments; LLaMA has shown competitive performance on many natural language processing (NLP) and IR tasks (Touvron et al., 2023b, a). Experiments on datasets from the TREC 2019–2022 deep learning (TREC-DL) tracks (Craswell et al., 2022, 2021, 2020, 2019) show that QPP-GenRE achieves state-of-the-art QPP quality in estimating the retrieval quality of a lexical ranker (BM25) and a neural ranker (ANCE, Xiong et al., 2021) in terms of RR@10, a precision-oriented IR metric, and nDCG@10, an IR metric considering recall. See Sections 6.1 and 6.2.
@@
We also find that using LLMs to directly model QPP, i.e., asking LLMs to directly generate values of IR evaluation metrics, performs much worse than QPP-GenRE. This finding reveals that QPP-GenRE is a more effective way of modeling QPP using LLMs. Furthermore, our experiments demonstrate the effectiveness of our devised approximation strategy in nDCG@10: QPP-GenRE achieves state-of-the-art QPP quality at the shallow judging depth 10, and QPP-GenRE’s QPP quality reaches saturation when it further judges up to 100–200 retrieved items in a ranked list. See Section 7.1.
Moreover, we investigate the impact of fine-tuning and the choice of LLMs on the quality of generated relevance judgments and QPP. Besides LLaMA-7B, we consider the newly-released Llama-3-8B (AI@Meta, 2024), and Llama-3-8B-Instruct (AI@Meta, 2024), and compare their fine-tuned versions against their few-shot prompted counterparts. We find that fine-tuning markedly improves the quality of relevance judgment generation and QPP for all LLMs; the performance of fine-tuned LLMs in terms of judging relevance exceeds that of a commercial LLM (GPT-3.5) (Faggioli et al., 2023a). Also, we observe that the newly released and instruction-tuned LLMs generally deliver better performance. This implies that with a more effective LLM, QPP-GenRE has the potential to be more effective at QPP. See Section 7.2.
Additionally, to show QPP-GenRE’s compatibility with other types of relevance prediction methods, we adapt a state-of-the-art pointwise LLM-based re-ranker, RankLLaMA (Ma et al., 2023a), into a relevance judgment generator by applying a threshold to its re-ranking scores. Our results indicate that QPP-GenRE integrated with RepLLaMA achieves high QPP quality, at the cost of tuning a proper threshold. The high QPP quality achieved by RankLLaMA demonstrates QPP-GenRE’s compatibility with other types of relevance prediction methods. See Section 7.3.
We also analyze QPP errors based on automatically generated relevance judgments, demonstrating QPP-GenRE’s interpretability. See Section 7.4.
Finally, our computational cost analysis shows that QPP-GenRE shows lower latency than some supervised QPP baselines when predicting multiple measures because multiple measures can be derived from the same set of relevance judgments. Although QPP-GenRE shows higher latency than other QPP baselines when predicting only one metric, QPP-GenRE ’s latency is still 20 times smaller than the state-of-the-art GPT-4-based listwise re-ranker (Sun et al., 2023b). See Section 7.5.
Application scenarios
@@
Given QPP-GenRE’s high QPP quality and interpretability, it is well-suited for some knowledge-intensive professional search scenarios, e.g., legal (Tomlinson et al., 2007) or patent search (Lupu et al., 2013), where accurate QPP is prioritized, interpretable QPP results are preferred, and users may have a higher tolerance level for latency than users in web search. Plus, QPP-GenRE can be used to analyze how well a search system performs in offline settings (Faggioli et al., 2023d), where latency is not necessarily an issue.
One might argue, if QPP-GenRE needs to be integrated with an LLM to predict ranking quality, why not directly use the LLM for re-ranking? However, we reveal that QPP-GenRE integrated with LLaMA-7B already achieves high QPP quality and remains significantly more efficient than costly state-of-the-art LLM-based re-rankers (e.g., the GPT-4-based listwise re-ranker (Sun et al., 2023b)). Calling those expensive LLM-based re-rankers is often unnecessary, as many initial rankings are good enough and either do not require re-ranking or only need very shallow re-ranking depths (Meng et al., 2024). Therefore, sufficiently accurate QPP for initial rankings is needed to guide the decision on whether to use the expensive re-ranker, or to determine the optimal re-ranking depth that does not waste computational resources. Given QPP-GenRE’s substantial improvements in QPP quality over previous QPP methods and significantly lower latency compared to those expensive re-rankers, it is valuable to make QPP-GenRE work with state-of-the-art, yet much more costly, LLM-based re-rankers (Sun et al., 2023b) to achieve a better balance between effectiveness and efficiency in re-ranking.
Reproducibility
To facilitate future research, we release our data, scripts for fine-tuning/inference, sampled demonstration examples for few-shot prompting, and fine-tuned checkpoints for LLaMA-7B, Llama-3-8B, and Llama-3-8B-Instruct at https://github.com/ChuanMeng/QPP-GenRE.
Contributions
Our main contributions are as follows:
@@
We propose a novel QPP framework using automatically generated relevance judgments (QPP-GenRE), which decomposes QPP into independent subtasks of predicting the relevance of each item in a ranked list to the query, and predicts different IR evaluation measures based on the relevance predictions.
@@
We devise an approximation strategy to predict IR measures considering recall, avoiding the cost of traversing the entire corpus to identify all relevant items for a query.
@@
We fine-tune leading open-source LLMs, LLaMA-7B, Llama-3-8B, and Llama-3-8B-Instruct for the task of automatically generating relevance judgments.
@@
We conduct experiments on four datasets, showing that QPP-GenRE outperforms the state-of-the-art QPP baselines on the TREC-DL 19–22 datasets in predicting RR@10 and nDCG@10 in terms of Pearson’s and Kendall’s .
@@ 2. Related Work
Our work is relevant to four strands of research: query performance prediction (QPP) (Section 2.1), zero/few-shot prompting and parameter-efficient fine-tuning (PEFT) for LLMs (Section 2.2), LLMs for generating relevance judgments (Section 2.3), and LLMs for re-ranking (Section 2.4).
2.1. Query performance prediction
QPP has attracted lots of attention in the IR and NLP community and has been widely studied in ad-hoc search (Singh et al., 2023; Faggioli et al., 2023e, f; Datta et al., 2022c), conversational search (Faggioli et al., 2023c, d; Abbasiantaeb et al., 2023; Meng et al., 2023c, a; Sun et al., 2021; Meng, 2024; Meng et al., 2023b), question answering (Samadi and Rafiei, 2023; Hashemi et al., 2019), and image retrieval (Poesina et al., 2023). This paper focuses on QPP for ad-hoc search.
A large number of unsupervised and supervised post-retrieval QPP methods have been proposed (Carmel and Yom-Tov, 2010) for predicting the performance of lexical rankers, such as query likelihood (Lafferty and Zhai, 2001) and BM25 (Robertson et al., 2009). Unsupervised QPP methods can be classified into clarity-based (Cronen-Townsend et al., 2002), robustness-based (Zhou and Croft, 2007, 2006; Aslam and Pavlu, 2007), coherence-based (Arabzadeh et al., 2021a; Diaz, 2007), and score-based (Tao and Wu, 2014; Shtok et al., 2012; Cummins et al., 2011; Pérez-Iglesias and Araujo, 2010; Zhou and Croft, 2007). More recently, a set of supervised QPP methods have been proposed (Arabzadeh et al., 2021b; Hashemi et al., 2019; Zamani et al., 2018; Datta et al., 2022a, b; Chen et al., 2022; Khodabakhsh and Bagheri, 2023). NeuralQPP (Zamani et al., 2018) and Deep-QPP (Datta et al., 2022a) are optimized from scratch. NQA-QPP (Hashemi et al., 2019) and BERT-QPP (Arabzadeh et al., 2021b) fine-tune BERT (Devlin et al., 2019) to improve QPP effectiveness. Further, Datta et al. (2022c) propose qppBERT-PL, which considers list-wise-document information, while Chen et al. (2022) propose BERT-groupwise-QPP that considers both cross-query and cross-document information. Khodabakhsh and Bagheri (2023) propose a multi-task query performance prediction framework (M-QPPF), learning document ranking and QPP simultaneously.
Post-retrieval QPP methods designed for lexical rankers struggle to predict the retrieval quality of neural rankers (Faggioli et al., 2023f; Hashemi et al., 2019), motivating several new unsupervised post-retrieval QPP methods designed for neural rankers. Datta et al. (2022b) propose a weighted relative information gain-based model (WRIG), which assesses a neural ranker for a given query by considering the relative difference of predicted performance between the given query and its variants; Zendel et al. (2023) assess a neural re-ranker by measuring the entropy of scores returned by it; Faggioli et al. (2023e) propose neural-ranker-specific ways of calculating regularization terms used by unsupervised post-retrieval QPP methods; Vlachou and Macdonald (2023) propose an unsupervised coherence-based QPP method that employs neural embedding representations to assess dense retrievers; and Singh et al. (2023) propose pairwise rank preference-based QPP (QPP-PRP) for predicting the performance of a neural ranker by measuring the degree to which a pairwise neural re-ranker (e.g., DuoT5 (Pradeep et al., 2021)) agrees with the ranked list returned by the neural ranker.
We present a novel QPP perspective: we start by automatically generating relevance judgments for a ranked list for a query and then proceed to predict IR evaluation measures for the ranked list. To the best of our knowledge, no prior work addresses QPP from this perspective.
Unlike regression-based QPP models (Khodabakhsh and Bagheri, 2023; Chen et al., 2022; Datta et al., 2022a; Arabzadeh et al., 2021b; Hashemi et al., 2019), which require training separate models to predict different IR evaluation measures, the training of LLMs for judging relevance in the QPP-GenRE method that we propose is agnostic to a specific IR evaluation measure, and different measures can be derived from the same set of generated relevance judgments.
We also differ from qppBERT-PL (Datta et al., 2022c), which first predicts the number of relevant items for each chunk in a ranked list and then aggregates those numbers into a general QPP score. However, qppBERT-PL’s output is still presented as a single scalar, which is insufficient to accurately represent different evaluation measures; also, it is infeasible to predict arbitrary IR measures only using the number of relevant items in a ranked list.
The work closest to QPP-GenRE, which is still different, is QPP using effectiveness evaluation without relevance judgments (EEwRJ) (Mizzaro et al., 2018). The goal of EEwRJ methods is to predict search system effectiveness in a TREC-like environment. E.g., a method proposed by Soboroff et al. (2001) randomly samples items from a pool for a query and treats these items as relevant; the intuition is that if an item is ranked highly by many search systems, it is likely to be pooled and therefore considered relevant. Mizzaro et al. (2018) explore applying QPP EEwRJ (Mizzaro et al., 2018) methods to QPP. However, QPP using EEwRJ suffers from two limitations: (i) EEwRJrequires obtaining ranked lists returned by all search systems in a given TREC edition to predict the difficulty of a query, and (ii) EEwRJencounters normalization challenges when predicting the ranking quality for a ranked list returned by a specific search system (Mizzaro et al., 2018). QPP-GenRE does not face these limitations.
2.2. Zero/few-shot prompting and parameter-efficient fine-tuning (PEFT) for LLMs
While fine-tuning pre-trained language models has given rise to many state-of-the-art results (Devlin et al., 2019), fully fine-tuning LLMs for a specific task on consumer-level hardware is typically infeasible (Zhu et al., 2023) because of the large number of parameters of LLMs. As a result, there are three prevailing ways to adapt LLMs to a specific task: zero-shot prompting, few-shot prompting, a.k.a. in-context learning (ICL) (Dong et al., 2022; Brown et al., 2020), and parameter-efficient fine-tuning (PEFT) (Dettmers et al., 2023; Liu et al., 2022; Hu et al., 2021).
There is limited success in only prompting open-source LLMs for certain tasks (Qin et al., 2023). Zero-shot prompting instructs an LLM to perform a specific task by inputting a text instruction. To get a promising result, zero-shot prompting is usually based on instruction-tuned LLMs (Zhang et al., 2023b; Qin et al., 2023), such as Flan-T5 (Chung et al., 2022), Flan-UL2 (Tay et al., 2022). However, Sun et al. (2023a) show that the performance of zero-shot prompting degrades considerably if an LLM is fed an instruction that was not observable during its training. ICL inputs a few input-target pairs (a.k.a. demonstrations) to an LLM, which would make an LLM learn from analogy (Dong et al., 2022) without updating its parameters. However, ICL has a high computational cost because it needs to feed input-target pairs to an LLM for each prediction; also, ICL requires substantial manual prompt engineering because an LLM’s performance (Liu et al., 2022) is sensitive to the formatting of the prompt (e.g., the wording and the order of input-target pairs).